OpenStudy (anonymous):

The largest box of popcorn in the USA measured 40ft long, 28 ft wide, and 6.7 ft high. It was filled with popcorn at Pittsville Elemantary in Wisconsin. How much popcorn filled the box ?

OpenStudy (anonymous):

 The box is in the shape of a rectangular prism. You are asked to find the volume of the box. The formula for the volume of a rectangular prism is\[V = lwh\]The dimensions are all given in the question. Your answer will be in \(\text{ft}^3\), or cubic feet.

OpenStudy (trey..):

121

OpenStudy (anonymous):

@ospreytriple I got 750.4

OpenStudy (anonymous):

@halia according to the dimensions given, your decimal is in the wrong place.

OpenStudy (anonymous):

75.04???

OpenStudy (anonymous):

Don't guess. Calculate \(40 \times 28 \times 6.7\)

OpenStudy (anonymous):

I did it on my calculator and it showed up as 7504

OpenStudy (anonymous):

Correct. Well done.

OpenStudy (anonymous):

So that's the volume ?

OpenStudy (anonymous):

@ospreytriple now im asked to find the surface area can you help me out ?

OpenStudy (anonymous):

Sure. How many surfaces does the box have?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Great. What shape are the surfaces?

OpenStudy (anonymous):

@ospreytriple they are squares and rectangles

OpenStudy (anonymous):

Well, I don't think there are any squares, but they are 6 rectangles. Do you know how to calculate the area of a rectangle?

OpenStudy (anonymous):

Um is it A=l*w*h ?

OpenStudy (anonymous):

Not quite. Area of a rectangle is \(A=lw\). Now with your box, keep in mind that each pair of opposite surfaces are the same size, i.e. the front and back are the same, the left side and right side are the same, and the top and bottom are the same. Understand?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

OK. Using what's given and referring to the sketch, the front and back each have an area of \(l \times h\), the left and right sides each have an area of \(w \times h\), and the top and bottom each have an area of \(l \times w\). Putting all the sides together we get, for the surface area of a rectangular prism,\[SA = 2lw + 2lh + 2wh\]Plug in the values and solve.

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

@ospreytriple I plugged them in and I got 2240+536+375.2

OpenStudy (anonymous):

Add them up and what do you get?

OpenStudy (anonymous):

351.2 ft

OpenStudy (anonymous):

No. Try again.

OpenStudy (anonymous):

3151.2 ft squared

OpenStudy (anonymous):

Excellent. Good job. That's your answer.

OpenStudy (anonymous):

Thank you so much !!!!!!! @ospreytriple I really appreciated :)

OpenStudy (anonymous):

You're welcome